solve for x
log3 + log(x+4) = log5 + log(x-3)

Answer by lwsshak3(6522)

solve for x
log3 + log(x+4) = log5 + log(x-3)
**
log3 + log(x+4) = log5 + log(x-3)
log3 + log(x+4) - log5 - log(x-3)=0
log3 + log(x+4) - (log5 + log(x-3))=0
place under a single log
log(3*(x+4))/(5*(x-3)=0
convert to exponential form: (base(10) raised to log of number(0)=number(3*(x+4))/(5*(x-3)
10^0=3x+12/5x-15=1
3x+12=5x-15
2x=27
x=27/2=13.5
Check:
log3 + log(x+4) = log5 + log(x-3)
log3 + log(17.5) = log5 + log(10.5)
1.7202=1.7202